# Trustworthy Pedestrian Trajectory Prediction via Pattern-Aware Interaction Modeling

### v2

@@ Abstract
Accurate and reliable pedestrian trajectory prediction is critical for the safety and robustness of intelligent applications, yet achieving trustworthy prediction remains highly challenging due to the complexity of interactions among pedestrians. Previous methods often adopt black-box modeling of pedestrian interactions, treating all neighbors uniformly. Despite their strong performance, such opaque modeling limits the reliability of predictions in safety-critical real-world deployments. To address this issue, we propose InSyn (Interaction-Synchronization Network), a novel Transformer-based model that explicitly captures diverse interaction patterns (e.g., walking in sync or conflicting) while effectively modeling direction-sensitive social behaviors. Additionally, we introduce a training strategy, termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction. Experiments on the ETH and UCY datasets demonstrate that our model not only outperforms recent black-box baselines in prediction accuracy, especially under high-density scenarios, but also provides stronger interpretability, achieving a favorable trade-off between reliability and accuracy. Furthermore, the SSOS strategy proves to be effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%.
Index Terms:
Human Trajectory Prediction, Interpretable Modeling, Attention Mechanism
I Introduction
@@
Pedestrian trajectory prediction is essential for safety-critical applications such as autonomous driving [1] and robotic navigation [2]. Accurate and reliable forecasting enables intelligent systems to understand human behavior better and ensures safe and trustworthy predictions. However, the presence of complex pedestrian interactions poses significant challenges to the task. In a given scene, pedestrians may respond to nearby individuals in different ways. For instance, the trajectories of two pedestrians may exhibit conflict. Alternatively, they may show weak or no influence on each other, such as when walking in sync or as part of a group. In particular, even in walk-in-sync scenarios, subtle interactions may still occur. Effectively modeling these intricate interaction patterns, especially in high-density environments, remains a major challenge in pedestrian trajectory prediction.
@@
Figure 1: Comparison of Interaction Modeling: Previous Methods vs. Our Approach. F represents the interaction effect between pedestrians. In traditional approaches (top), all neighbors of the agent are treated as being in the same state, with interactions solely based on the distance between the agent and its neighbors. This simplistic assumption may lead to overestimating the influence of fellow pedestrians. Our method (bottom) introduces a more refined modeling strategy by considering the specific states of neighboring pedestrians (e.g., In Sync or Conflict), providing a more nuanced understanding of the neighboring interaction.
@@
Recently, extensive studies have investigated pedestrian interactions [3, 4, 5, 6], using approaches like social pooling layers [3, 7], social force mechanism [6], and graph neural networks (GNNs) [8, 9]. However, a common characteristic of these methods is that they rely on relative positions between an agent and its neighbors to model influence, treat all neighboring pedestrians by black-box representation, failing to differentiate patterns such as In Sync and Conflict (see Figure 1). This oversimplification may lead to overfitting and untrustworthy predictions. For example, ignoring synchronized walking (e.g., friends walking side-by-side) can lead to overlapping predicted paths. In conflict cases, such as head-on encounters, the absence of explicit modeling may cause unrealistic predictions that either overestimate collisions or miss necessary evasive actions. This black-box modeling not only reduces prediction accuracy, but also limits interpretability and trustworthiness in safety-critical scenarios such as machine navigation and intelligent transportation systems. These limitations become particularly pronounced in crowded or dynamic environments.
@@
To address this issue, we need a method that understands what’s exactly going on between pedestrians—recognizing concrete interaction patterns rather than treating them as abstract black-box inputs—to ensure more reliable and interpretable predictions (see Figure 1).
@@
Although pedestrian trajectories are highly stochastic, they usually have a clear goal [10]. By focusing on the target goal, the performance of the model can be significantly enhanced. Experimental results from Giuliari et al. [11] demonstrate that the attention mechanism is well-suited for processing sequential data with irregular time intervals (e.g., observed trajectories and goals), effectively capturing long-term dependencies in pedestrian motion and providing reliable goal-driven functionality.
@@
This paper proposes the InSyn, a Transformer-based model. The model consists of three components: (1) Interaction Encoder: designed to explicitly extract interaction information and integrates interaction features, goal-driven behavior, and observed trajectories through the self-attention mechanism. (2) Trajectory Generator: forms an attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step. (3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goal sampling.
We evaluate our model on the well-established pedestrian trajectory prediction datasets ETH [12] and UCY [13]. The experimental results demonstrate that our model outperforms recent methods in key metrics, particularly in scenarios with complex interactions. We further conduct ablation studies to show the superiority of interaction modeling and the SSOS strategy. Additionally, the case study highlights the interpretability of our approach by illustrating how specific interaction patterns influence prediction outcomes.
Figure 2: Left is the illustration of the input $S_{0:\tau}$. At each time step, the walking state $S_{k}$ comprises the 2D coordinates $(x,y)$ and interaction information $N_{k}$. Right demonstrates the scenarios of In Sync, Conflict and No Neighbor state.
@@
II Related Work
II-A Pedestrian Trajectory Prediction
Pedestrian trajectory prediction methods can be categorized into traditional machine learning approaches and deep learning-based approaches. Traditional methods, such as Markov decision processes [14] and Gaussian distributions [15], are efficient in handling simple scenarios. However, they often struggle with complex data distributions and exhibit limited generalization capabilities. In contrast, deep learning approaches have shown significant improvements in modeling complex patterns [16, 17, 18, 19]. Long Short-Term Memory (LSTM) network [20], a classic time-series model, is applied in this field. Social-LSTM [3] introduced a pooling mechanism to incorporate interactions into trajectory prediction. Yang et al. [21] proposed a variant LSTM structure for information sharing within scenes, while SR-LSTM [22] combined neighbor intentions and message passing to enhance prediction accuracy. These works leverage the strong temporal feature extraction capabilities of LSTM.
In recent years, Transformers have gained popularity in trajectory prediction due to their ability to capture long-range dependencies. Liu et al. [23] combined CNN and Transformer, using the attention mechanism to process interaction information. MRGTraj [24] integrated a Transformer encoder with a novel decoder, achieving trajectory prediction through a mapping-refining-generating structure. CITraNet [25] significantly improved prediction accuracy by introducing an innovative Transformer-based Gumbel distribution network.
Building on these advancements, our research combines the strengths of LSTM and Transformer. By leveraging the temporal feature extraction capabilities of LSTM, we enhance the Transformer’s attention mechanism with processed temporal interaction features, thereby improving prediction performance.
II-B Pedestrian Interaction Modeling
@@
Current mainstream methods for interaction modeling include social pooling layers [3, 26], GNNs [9, 8], and attention mechanisms [27, 28]. Among these, the social attention module aggregates interaction information by analyzing correlations between pedestrian motion and future trajectories [26]. Yang et al. [28] introduced a social graph attention mechanism combined with a pseudo-oracle predictor to capture social interactions and intention states, enhancing prediction accuracy. Transformers have been utilized in pedestrian interaction modeling [29, 30, 31, 32, 33]. For example, Yuan et al. [34] proposed AgentFormer that learns spatiotemporal interaction embeddings from sequential trajectory features. Similarly, Yang et al. [30] leveraged GNNs and Transformers to model spatial and temporal dependencies. To address future interaction modeling, Amirloo et al. [32] employed the self-attention mechanism of Transformer to model pedestrian interactions and considered future states autoregressively during decoding to avoid trajectory conflicts.
@@
However, most of these methods rely on black-box representations of social influence. This limits the model’s ability to produce reliable and interpretable predictions, especially in complex or safety-critical scenarios. In contrast, our approach explicitly captures structured interaction patterns, such as walking In Sync or engaging in Conflict, aiming to enhance the trustworthiness and transparency of socially-aware trajectory prediction.
III Problem Formulation
@@
In this study, pedestrian trajectory prediction is formulated as follows: given the walking states $S_{0:\tau}$ of a certain agent during the observation time steps $\mathord{0:\tau}$, the model outputs the predicted trajectory $\widehat{pos}_{\tau+1:T}$ of the agent for the future time steps $\mathord{\tau+1:T}$. The input walking states of agent $p$ are denoted as $\mathcal{S}^{p}=\left(S_{0}^{p},S_{1}^{p},\dots,S_{\tau}^{p}\right)$, where $S_{t}^{p}$ represents the walking state at time step $t\ (0\leq t\leq\tau)$. Each walking state $S_{t}^{p}$ consists of the agent’s position $pos_{t}=(x_{t},y_{t})$ in 2D space and the pedestrian interaction information $N_{t}$. Since interactions are directional, we adopt a region partition strategy around each agent. The interaction information $N_{t}$ is defined as
$N_{t}=\left[\left(f_{t}^{lu},d_{t}^{lu}\right),\left(f_{t}^{ru},d_{t}^{ru}\right),\left(f_{t}^{ld},d_{t}^{ld}\right),\left(f_{t}^{rd},d_{t}^{rd}\right)\right]$ (1)
where $f_{t}^{lu},f_{t}^{ru},f_{t}^{ld},f_{t}^{rd}$ represent the interaction states in the left-up, right-up, left-down, and right-down regions, respectively. These patterns include No Neighbor, In Sync, and Conflict, as shown in Figure 2. The construction of interaction patterns is highly transparent and trustworthy, as it relies on interpretable spatial-temporal neighbor changes.
@@
Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time t1t-1, the interaction state is classified as In Sync; otherwise, it is marked as Conflict. In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be negligible. However, if a new pedestrian suddenly enters a region, referred to as Conflict, it is more likely to impact the agent’s trajectory. The terms $d_{t}^{lu},d_{t}^{ru},d_{t}^{ld},d_{t}^{rd}$ indicate the distances to the nearest neighbor in each region. A region with no pedestrians is classified as No Neighbor, and a large distance value is assigned to indicate minimal influence. The goal of this study is to train a generative model
$m_{\theta}\left(\widehat{pos}_{\tau+1:T}\mid S_{0:\tau}\right)$ (2)
@@
where $\theta$ represents the model parameters. The model aims to predict the distribution of pedestrians’ future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
@@ IV Methodology
We first discuss how our model captures specific interactions and achieves goal-driven functionality. We then elaborate on the motivation behind the SSOS strategy and, finally, the design of the Seq-CVAE module.
IV-A Interaction Encoder
@@
IV-A1 Goal-Driven
Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  IV-B) is combined with the position information through concatenation. Notably, since self-attention cannot inherently capture positional information, positional encoding is used to encode the input positions. This encoding is based on the sequence positions of the inputs. Therefore, before concatenating $pos_{0:\tau}^{p}$ with the sampled goal $\widehat{pos}_{T}$, padding needs to be inserted between them to ensure that the goal is assigned the corresponding positional encoding.
IV-A2 Interaction
To facilitate trustworthy modeling of social interactions, we introduce a modular neighbor encoder that encodes pattern-aware and distance-based interaction effects.
The interaction patterns between pedestrians are extracted through a neighbor encoder module followed by an LSTM. The agent $p$’s social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder. which consists of three components: an embedding layer, a Distance Gate, and a Social Pooling Layer. The embedding layer maps interaction states (No Neighbor, In Sync, Conflict) to learnable representations. These representations are then combined with the processed distance $d$ through element-wise multiplication. The process of $d$ is through the Distance Gate: $\mathrm{Gate}(d)=\sigma(\mathbf{W}\cdot d+\mathbf{b})$, where $\sigma(z)=\frac{1}{1+e^{-z}}$ and $\mathbf{W},\mathbf{b}$ represent the parameters of linear layer. Inspired by the gating mechanism in LSTM [20], the Distance Gate maps $d$ to a weight in [0,1], reflecting distance-based interaction intensity.
@@ IV-A2 Interaction
IV-B SSOS: Seq-Start of Seq
[…]
Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction—a numerical time-series prediction task—using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$. That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the ground truth $pos_{\tau+1}$. This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with the SOS and the cross-attention with the encoder’s output. Insufficient information or excessive noise in the SOS may lead to alignment bias, thus reducing prediction accuracy.
To mitigate this issue, we propose using the sequence data $pos_{0,1,...,\tau}$ as the SOS. We refer to this approach as SSOS and use it during training. In addition to computing the loss of the predicted trajectory $\widehat{pos}_{\tau+1:T-1}$, the loss of the reconstructed observed trajectory $\widehat{pos}_{1:\tau}$ is also minimized through gradient descent. The loss function using SSOS is presented below:

### v3

@@ Abstract
Accurate and reliable pedestrian trajectory prediction is critical for the application of intelligent applications, yet achieving trustworthy prediction remains highly challenging due to the complexity of interactions among pedestrians. Previous methods often adopt black-box modeling of pedestrian interactions. Despite their strong performance, such opaque modeling limits the reliability of predictions in real-world deployments. To address this issue, we propose InSyn (Interaction-Synchronization Network), a novel Transformer-based model that explicitly captures diverse interaction patterns (e.g., walking in sync or conflicting) while effectively modeling direction-sensitive social behaviors. Additionally, we introduce a training strategy, termed Seq-Start of Seq (SSOS), designed to alleviate the common issue of initial-step divergence in numerical time-series prediction. Experiments on the ETH and UCY datasets demonstrate that our model not only outperforms recent black-box baselines in prediction accuracy, especially under high-density scenarios, but also provides transparent interaction modeling, as shown in the case study. Furthermore, the SSOS strategy proves to be effective in improving sequential prediction performance, reducing the initial-step prediction error by approximately 6.58%. Code is avaliable at https://github.com/rickzky1001/InSyn
1 Introduction
@@
Pedestrian trajectory prediction is essential for real-world applications such as autonomous driving [18] and robotic navigation [3]. Accurate and reliable forecasting enables intelligent systems to understand human behavior better and ensures trustworthy predictions. However, the presence of complex pedestrian interactions poses significant challenges to the task. In a given scene, pedestrians may respond to nearby individuals in different ways [13]. For instance, the trajectories of two pedestrians may exhibit conflict. Alternatively, they may show weak or no influence on each other, such as when walking in sync or as part of a group. In particular, even in walk-in-sync scenarios, subtle interactions may still occur. Effectively modeling these intricate interaction patterns, especially in high-density environments, remains a major challenge in pedestrian trajectory prediction.
@@
Figure 1: Comparison of Interaction Modeling: Previous Methods vs. Our Approach. F represents the interaction effect between pedestrians. In traditional approaches (top), all neighbors of the agent are treated as being in the same state. Our method (bottom) introduces a more refined modeling strategy by considering the specific states of neighboring pedestrians, providing a more nuanced understanding of the neighboring interaction.
@@
Recently, extensive studies have investigated pedestrian interactions [1, 31, 23, 19, 40], using approaches like social pooling layers [1, 6], social force mechanism [40, 41], and graph neural networks (GNNs) [38, 17, 4]. However, a common characteristic of these methods is that they rely on features like relative position between an agent and its neighbors to model influence, treat all neighboring pedestrians by black-box representation, failing to differentiate patterns such as In Sync and Conflict (see Figure 1). This oversimplification may lead to overfitting and untrustworthy predictions. For example, failing to recognize synchronized walking (e.g., friends walking side-by-side) and mistakenly modeling them with repulsive forces can lead to unrealistic divergence in their predicted paths. This black-box modeling not only reduces prediction accuracy, but also limits interpretability and trustworthiness in real-world scenarios. These limitations become particularly pronounced in crowded or dynamic environments.
@@
To address this issue, we need a method that understands what’s exactly going on between pedestrians—recognizing interaction patterns rather than treating them as abstract black-box inputs—to ensure more reliable and accurate predictions (see Figure 1).
@@
Although pedestrian trajectories are highly stochastic, they usually have a clear goal [27]. Focusing on the target goal can significantly enhance model performance. Building on this insight and the suitability of the attention mechanism for capturing long-range dependencies in sequential data [7], we propose InSyn, a Transformer-based model.
@@
The model consists of three components: (1) Interaction Encoder: designed to explicitly extract interaction information and integrates interaction features, goal-driven behavior, and observed trajectories through the self-attention mechanism. (2) Trajectory Generator: forms an cross-attention mechanism in conjunction with the Interaction Encoder, and incorporates the proposed SSOS strategy to alleviate divergence at the initial prediction step. (3) Seq-CVAE Goal Sampler: a conditional generative model specifically designed for sequential prediction, used for goal sampling.
@@
2 Related Work
@@
Current mainstream methods for interaction modeling include social pooling layers [1, 34, 10], GNNs [17, 38, 24, 29], and attention mechanisms [16, 35, 20]. Among these, the social attention module aggregates interaction information by analyzing correlations between pedestrian motion and future trajectories [34]. Yang et al. [35] introduced a social graph attention mechanism combined with a pseudo-oracle predictor to capture social interactions and intention states, enhancing prediction accuracy. Transformers have been utilized in pedestrian interaction modeling [33, 36, 11, 2, 26]. For example, Yuan et al. [39] proposed AgentFormer that learns spatiotemporal interaction embeddings from sequential trajectory features. Similarly, Yang et al. [36] leveraged GNNs and Transformers to model spatial and temporal dependencies. To address future interaction modeling, Amirloo et al. [2] employed the self-attention mechanism of Transformer to model pedestrian interactions and considered future states autoregressively during decoding to avoid trajectory conflicts.
@@
However, most of these methods rely on black-box representations of social influence. This limits the model’s ability to produce reliable and interpretable predictions. In contrast, our approach explicitly leverages structured interaction patterns, such as walking In Sync or engaging in Conflict, aiming to enhance the trustworthiness and transparency of socially-aware trajectory prediction.
3 Problem Formulation
Figure 2: Left is the illustration of the input $S_{0:\tau}$. At each time step, the walking state $S_{k}$ comprises the 2D coordinates $(x,y)$ and interaction information $N_{k}$. Right demonstrates the scenarios of In Sync, Conflict and No Neighbor state. Note that the 4-regions partition and the scenarios identification use simple spatial-temporal rules for transparency, more complex methods could be applied here but beyond our scope.
@@ 3 Problem Formulation
In this study, pedestrian trajectory prediction is formulated as follows: given the walking states $S_{0:\tau}$ of a certain agent during the observation time steps $\mathord{0:\tau}$, the model outputs the predicted trajectory $\widehat{pos}_{\tau+1:T}$ of the agent for the future time steps $\mathord{\tau+1:T}$. The input walking states of agent $p$ are denoted as $\mathcal{S}^{p}=\left(S_{0}^{p},S_{1}^{p},\dots,S_{\tau}^{p}\right)$, where $S_{t}^{p}$ represents the walking state at time step $t\ (0\leq t\leq\tau)$. Each walking state $S_{t}^{p}$ consists of the agent’s position $pos_{t}=(x_{t},y_{t})$ in 2D space and the pedestrian interaction information $N_{t}$. Since interactions are directional, we adopt a region partition strategy [1] around each agent. The interaction information $N_{t}$ is defined as
$N_{t}=\left\{(f_{t}^{r},d_{t}^{r})\right\}_{r\in\{\mathrm{lu,ru,ld,rd}\}}$ (1)
where $f_{t}^{r}$ represents the interaction state in region $r$ at time $t$; $d_{t}^{r}$ indicates the distances to the nearest neighbor; $\{\mathrm{lu,ru,ld,rd}\}$ represent the left-up, right-up, left-down, right-down of the agent respectively.
@@
The interaction state $f_{t}^{r}$ can be No Neighbor, In Sync, and Conflict, as shown in Figure 2. The construction of interaction patterns is highly transparent. Specifically, if the nearest neighbor in a given region at time $t$ is the same as the one at time t1t-1, the interaction state is classified as In Sync; otherwise, it is marked as Conflict. A region with no pedestrians is classified as No Neighbor, and a large distance value is assigned to indicate minimal influence. In cases of No Neighbor or In Sync, the interaction influence of the neighbor on the agent’s future trajectory may be limited. However, if a new pedestrian suddenly enters a region, referred to as Conflict, it is more likely to impact the agent’s trajectory. This identification uses simple rules for transparency and it could be replaced by a model-based method.
@@
The goal of this study is to train a generative model $m_{\theta}\left(\widehat{pos}_{\tau+1:T}\mid S_{0:\tau}\right)$, aiming to predict the distribution of pedestrians’ future trajectory over time steps $\mathord{\tau+1:T}$, conditioned on the walking states during the observation period $\mathord{0:\tau}$.
4 Methodology
@@
4.1 Interaction Encoder
@@
4.1.1 Goal-Driven.
Given the history position information $pos_{0:\tau}^{p}$, the goal $\widehat{pos}_{T}$ generated by Seq-CVAE (introduced in Section  4.3) is combined with the position information through concatenation. Notably, since self-attention cannot inherently capture positional information, positional encoding is used to encode the input positions. This encoding is based on the sequence positions of the inputs. Therefore, to ensure the sampled goal $\widehat{pos}_{T}$ is assigned the correct positional encoding, a $T-\tau-1$ length padding tokens must be inserted between the observed trajectory $pos{0:\tau}^{p}$ and the goal before concatenation.
4.1.2 Interaction.
To facilitate trustworthy modeling of social interactions, we introduce a modular neighbor encoder that encodes pattern-aware effects.
The interaction patterns between pedestrians are extracted through a neighbor encoder module followed by an LSTM. The agent $p$’s observed social interaction information, represented as $\mathcal{N}^{p}=\left(N_{0}^{p},N_{1}^{p},...,N_{\tau}^{p}\right)$, are encoded using the neighbor encoder. which consists of three components: an embedding layer, a Distance Gate, and a Social Pooling Layer, as shown in Figure 3. The embedding layer maps interaction states $f_{t}^{r}$ to learnable representations. These representations are then combined with the processed distance $d$ through element-wise multiplication. The process of $d$ is through the Distance Gate: $\mathrm{Gate}(d)=\sigma(\mathbf{W}\cdot d+\mathbf{b})$, where $\sigma(z)=\frac{1}{1+e^{-z}}$ and $\mathbf{W},\mathbf{b}$ represent the parameters of linear layer. Inspired by the gating mechanism in LSTM [9], the Distance Gate maps $d$ to a weight in [0,1], reflecting distance-based interaction intensity.
@@
4.2 Trajectory Generator
The Trajectory Generator incorporates the SSOS Strategy into the Transformer Decoder [28] to alleviate initial prediction divergence. The Transformer’s decoder relies on the Start of Sequence (SOS) as the initial input when generating the first output. In Natural Language Processing (NLP), the Begin-of-Sentence token $\left<bos\right$> is commonly used as the SOS, which aligns with the semantic structure of text. However, in sequential trajectory prediction, the selection of SOS should consider the spatial information; otherwise, it may introduce noise. Giuliari et al. [7] set the SOS to $(0,0)$ for trajectory prediction. While $(0,0)$ represents a fixed position in spatial coordinates, using it as the SOS may introduce noise and mislead the model due to its lack of alignment with the observed trajectory. Therefore, it is reasonable to consider using the position at the last observed time step, $\widehat{pos}_{\tau}$, as the SOS token of the decoder.
Nevertheless, we note that when the Transformer’s encoder-decoder architecture is applied to trajectory prediction, a numerical time-series prediction task, using a single value as the SOS can lead to a lack of smoothness in the transition between the first predicted value $\widehat{pos}_{\tau+1}$ and the observed trajectory $pos_{0:\tau}$. That is to say, $\widehat{pos}_{\tau+1}$ may deviate noticeably from the last observation point $pos_{\tau}$. This issue likely arises because, during the initial decoding stage, the Transformer decoder relies solely on the self-attention with the SOS and the cross-attention with the encoder’s output. Insufficient information or excessive noise in the SOS may lead to alignment bias, thus reducing prediction accuracy.
To mitigate this issue, we propose using the sequence data $pos_{0:\tau}$ as the SOS. We refer to this approach as SSOS and use it during training. In addition to computing the loss of the predicted trajectory $\widehat{pos}_{\tau+1:T-1}$, the loss of the reconstructed observed trajectory $\widehat{pos}_{1:\tau}$ is also minimized through gradient descent. The loss function using SSOS is presented below: